OpenStudy (anonymous):

Write the function 2 f (x ) = -2x^2 - 20x + 15

OpenStudy (anonymous):

I think it starts off as : 2 f (x ) = -2(x + 5)^2 but im not sure about k

OpenStudy (anonymous):

f(x)= is the same as saying y= so... 2f(x)= 2y= so... f(x)= blahblahblah ------------ ? 2

OpenStudy (anonymous):

nevermind the 2 in front of the fx but I have to write it in this form: f(x ) = a (x -h ) ^+ k

OpenStudy (anonymous):

should I use complete the square method but wouldn't it say to do that in the questions?

OpenStudy (anonymous):

I skipped algebra, what is the name of the form? not vertex form is it?

OpenStudy (anonymous):

standard form

OpenStudy (anonymous):

it is in standard form, you are just trying to factor it?

OpenStudy (anonymous):

i would use quadratic myself unless they said to complete the square

OpenStudy (anonymous):

the answer is -2(x+5)^2 and the last number K could be -65,35,65,115. I don't think its the quadratic formula

OpenStudy (anonymous):

that was the form i was asking about, with the k value

OpenStudy (anonymous):

that's : f (x ) = a (x -h )^2 + k

OpenStudy (anonymous):

vertex form?

OpenStudy (anonymous):

but no matter what I do I cant figure out K

OpenStudy (anonymous):

it only says standard form

OpenStudy (anonymous):

try k 65

OpenStudy (anonymous):

65-2 (x+5)^2

OpenStudy (anonymous):

okay Nm I do have to use completing the square .... thank you! =)

OpenStudy (anonymous):

yeah i prefer quadratic but they both work

OpenStudy (anonymous):

sorry, I took pre algebra years ago and when I went back to school I tested out past algebra 2 and geometry. I'm doing fine in calculus but there are some common terms I don't know that I have to look up.